OpenStudy (anonymous):

Find the value by referring to the graph of the sine function. sin(3/2pi)

OpenStudy (anonymous):

0.45952901044

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

you just gotta plug the number into the calculator

OpenStudy (anonymous):

oh wait hold on mis understood the question

OpenStudy (anonymous):

I don't think that's how I was supposed to do it.

OpenStudy (anonymous):

do you need like a number value because of you do then all you do is plug it in (3/2pi) pi =180 2pi=360

OpenStudy (anonymous):

\[\sin (\frac{ 3 }{ 2 }\pi)\] im supposed to use a graph

OpenStudy (anonymous):

The angle is in the 2nd quadrant and the angle of reference (always from the x axis) is 60 degrees. The sin of 60 is the square root of 3 over 2. And, since it's in the 2nd quadrant, it is positive. Remember sin is positive in the 1st and 2nd quadrant and negative in the 3rd and 4th. So you should be able to use your graph now Sources: Friend + Google (for clarification)

OpenStudy (anonymous):

i still have no idea what to do lol. The graph won't even show up on my graphing calculator

OpenStudy (anonymous):

What buttons are you pushing? .-. lol

OpenStudy (anonymous):

it just says error when i try to graph it lol

OpenStudy (anonymous):

I hate calculators -,- the equation and the work is right your calculator just hates us lol ;)

OpenStudy (anonymous):

i still dont know how to find my answer tho

OpenStudy (anonymous):

how do i explain it?

OpenStudy (anonymous):

like better i guess lol but hey if this doesnt work i know how to make the day a little bit better lol

OpenStudy (anonymous):

?

OpenStudy (anonymous):

http://www.youtube.com/watch?v=IT4GPj4cwpA come on that funny lol but back to work how to i explain it more?